# AllTracker: Efficient Dense Point Tracking at High Resolution

### v1

@@ Abstract
We introduce AllTracker: a model that estimates long-range point tracks by way of estimating the flow field between a query frame and every other frame of a video. Unlike existing point tracking methods, our approach delivers high-resolution and dense (all-pixel) correspondence fields, which can be visualized as flow maps. Unlike existing optical flow methods, our approach corresponds one frame to hundreds of subsequent frames, rather than just the next frame. We develop a new architecture for this task, blending techniques from existing work in optical flow and point tracking: the model performs iterative inference on low-resolution grids of correspondence estimates, propagating information spatially via 2D convolution layers, and propagating information temporally via pixel-aligned attention layers. The model is fast and parameter-efficient (16 million parameters), and delivers state-of-the-art point tracking accuracy at high resolution (i.e., tracking pixels, on a 40G GPU). A benefit of our design is that we can train on a wider set of datasets, and we find that doing so is crucial for top performance. We provide an extensive ablation study on our architecture details and training recipe, making it clear which details matter most. Our code and model weights are available: alltracker.github.io
@@ 2 Related Work
The current state-of-the-art in optical flow estimation is SEA-RAFT [54]. SEA-RAFT first directly estimates a low-resolution flow field [14, 22], then computes appearance features with a ResNet-34 [20], and computes pixel-to-pixel correlations with these features [58], then iteratively estimates refinements to the flow via 2D convolutions that take flows and correlations as input, and finally upsamples the flow to full resolution via a pixel-shuffle layer [41]. Our work takes inspiration from SEA-RAFT (and its predecessor RAFT [47]) in the way that we iteratively refine low-resolution flows and upsample to full resolution, but we operate on multiple frames at a time (i.e., 16 frames instead of 2), and we add temporal attention layers to help information propagate across the time axis.
@@
The current state-of-the-art in point tracking is CoTracker3 [25]. CoTracker3 first computes feature maps for all frames, then initializes tracks for a sparse set of query points, then retrieves local correlations based on these tracks, and interleaves modules which pass information within tracks (temporally) and across tracks (spatially). The main novelty is in the spatial propagation: each point attends to a set of latent “virtual points” which learn a compressed representation of the video motions. While CoTracker3 delivers very accurate tracks, it requires users to pre-select a sparse set of points to track, and results vary based on how the queries are distributed [23]. CoTracker3 is more efficient than its predecessors [23, 19], but is still limited to tracking a few thousand points at a time. Our approach is simpler and more memory-efficient: we track on a low-resolution grid, where we achieve message-passing with simple 2D convolutions, and we recover spatial precision at the end of the architecture with a fast upsampling layer [41]. We note that this efficient upsampling technique was known well before the recent resurgence of point trackers [41], but perhaps its relevance was not fully appreciated.
@@ 3.6 Implementation details
We provide additional implementation details in the supplementary material, but describe the main details here.
@@ 4 Experiments
This section summarizes the key results, which concern the accuracy of our tracker compared to prior state-of-the-art, on a diverse array of test benchmarks. We provide additional details for our experiments in the appendices.
@@ 4.1 Main results
As shown in the table, even our weakest variant, AllTracker-Tiny-Kub, is competitive with state-of-the-art on many datasets, despite being only 6.29M parameters and training on a single synthetic dataset. Our full AllTracker model outperforms all other models on average, with the closest competitor being the variant of CoTracker3 which relied on an expensive bootstrapping scheme (noted by “+15k” in the table). This CoTracker3 model wins on three datasets (CroHD, Davis, DriveTrack) and our model wins on the rest and wins on average (66.1 vs. 65.0). Our method has a substantial gain over past work on the RGB-Stacking benchmark, which has many query points inside textureless regions; our model’s good performance here suggests that it is able to incorporate spatial context from a wider area than the previous models. We note that CoTracker3’s bootstrapping technique could be combined with our contributions, but it is simpler to gather additional synthetic data as we have done.
@@ 5 Conclusion and Limitations
AllTracker is an approach to point tracking that treats the task as multi-frame optical flow. Our model delivers state-of-the-art performance on point tracking benchmarks, and produces dense output at high resolution, which past point trackers have struggled to do. AllTracker makes sparse point trackers (of similar speed and accuracy) redundant, but interestingly it does not make optical flow methods redundant: it does not outperform the state-of-the-art optical flow methods on optical flow estimation. The model appears to be underfitting on short-range motion estimation, suggesting that better models might be obtained with greater compute. A related limitation to address is the temporal window size: with bigger GPUs it may be possible to simply train and test with a wider windows [25], and resolve longer occlusions. A broader area for future work is to add awareness about physical and common-sense constraints on motion, perhaps using 3D [56, 37] or more expressive model designs [18].
@@ Appendix F Additional qualitative results
We show additional visualizations of multiple models’ dense outputs in Figure 6.
@@
We notice striking dissimilarity across the outputs of the methods, attesting to the difficulty of the task, and to the usefulness of visualizing point tracks as flow maps. We find that when the foreground displacements are large, the flow models often “give up” on the dynamic foreground and produce a motion field that only describes the background. We also find that PIPs++ and LocoTrack often struggle with spatial smoothness, while the flow models do not. CoTracker3 occasionally fails on smoothness too (see row 3 with the car), but less so. Our model appears to produce results that are smooth and accurate, which matches intuitions for a model that blends the 2D processing of flow models with the temporal coherence of point trackers.

### v2

@@ Abstract
We introduce AllTracker: a model that estimates long-range point tracks by way of estimating the flow field between a query frame and every other frame of a video. Unlike existing point tracking methods, our approach delivers high-resolution and dense (all-pixel) correspondence fields, which can be visualized as flow maps. Unlike existing optical flow methods, our approach corresponds one frame to hundreds of subsequent frames, rather than just the next frame. We develop a new architecture for this task, blending techniques from existing work in optical flow and point tracking: the model performs iterative inference on low-resolution grids of correspondence estimates, propagating information spatially via 2D convolution layers, and propagating information temporally via pixel-aligned attention layers. The model is fast and parameter-efficient (16 million parameters), and delivers state-of-the-art point tracking accuracy at high resolution (i.e., tracking pixels, on a 40G GPU). A benefit of our design is that we can train jointly on optical flow datasets and point tracking datasets, and we find that doing so is crucial for top performance. We provide an extensive ablation study on our architecture details and training recipe, making it clear which details matter most. Our code and model weights are available: https://alltracker.github.io
@@ 2 Related Work
The current state of the art in optical flow estimation is SEA-RAFT [54]. SEA-RAFT first directly estimates a low-resolution flow field [14, 22], then computes appearance features with a ResNet-34 [20], and computes pixel-to-pixel correlations with these features [58], then iteratively estimates refinements to the flow via 2D convolutions that take flows and correlations as input, and finally upsamples the flow to full resolution via a pixel-shuffle layer [41]. Our work takes inspiration from SEA-RAFT (and its predecessor RAFT [47]) in the way that we iteratively refine low-resolution flows and upsample to full resolution, but we operate on multiple frames at a time (i.e., 16 frames instead of 2), and we add temporal attention layers to help information propagate across the time axis.
@@
The current state of the art in point tracking is CoTracker3 [24]. CoTracker3 first computes feature maps for all frames, then initializes tracks for a sparse set of query points, then retrieves local correlations based on these tracks, and interleaves modules which pass information within tracks (temporally) and across tracks (spatially). The main novelty is in the spatial propagation: each point attends to a set of latent “virtual points” which learn a compressed representation of the video motions. While CoTracker3 delivers very accurate tracks, it requires users to pre-select a sparse set of points to track, and results vary based on how the queries are distributed [25]. CoTracker3 is more efficient than its predecessors [25, 19], but is still limited to tracking a few thousand points at a time. Our approach is simpler and more memory-efficient: we track on a low-resolution grid, where we achieve message-passing with simple 2D convolutions, and we recover spatial precision at the end of the architecture with a fast upsampling layer [41]. We note that this efficient upsampling technique was known well before the recent resurgence of point trackers [41], but perhaps its relevance was not fully appreciated.
@@ 3.6 Implementation details
We provide extensive implementation details in the supplementary material, but describe the main details here.
@@ 4 Experiments
This section summarizes the key results, which concern the accuracy of our tracker compared to prior state of the art, on a diverse array of test benchmarks. We provide additional details for our experiments in the appendices.
@@ 4.1 Main results
As shown in the table, even our weakest variant, AllTracker-Tiny-Kub, is competitive with state of the art on many datasets, despite being only 6.29M parameters and training on a single synthetic dataset. Our full AllTracker model outperforms all other models on average, with the closest competitor being the variant of CoTracker3 which relied on an expensive bootstrapping scheme (noted by “+15k” in the table). This CoTracker3 model wins on three datasets (CroHD, Davis, DriveTrack) and our model wins on the rest and wins on average (66.1 vs. 65.0). Our method has a substantial gain over past work on the RGB-Stacking benchmark, which has many query points inside textureless regions; our model’s good performance here suggests that it is able to incorporate spatial context from a wider area than the previous models. We note that CoTracker3’s bootstrapping technique could be combined with our contributions, but it is simpler to gather additional synthetic data as we have done.
@@ 5 Conclusion and Limitations
AllTracker is an approach to point tracking that treats the task as multi-frame optical flow. Our model delivers state-of-the-art performance on point tracking benchmarks, and produces dense output at high resolution, which past point trackers have struggled to do. AllTracker makes sparse point trackers (of similar speed and accuracy) redundant, but interestingly it does not make optical flow methods redundant: it does not outperform the state-of-the-art optical flow methods on optical flow estimation. The model appears to be underfitting on short-range motion estimation, suggesting that better models might be obtained with greater compute. A related limitation to address is the temporal window size: with bigger GPUs it may be possible to simply train and test with wider windows [24], and resolve longer occlusions. A broader area for future work is to add awareness about physical and common-sense constraints on motion, perhaps using 3D [56, 37] or more expressive model designs [18, 60].
@@ Appendix B Additional training details
We train with mixed precision in PyTorch (bfloat16).
@@ Appendix F Additional qualitative results
We show additional visualizations of multiple models’ dense outputs in Figure 6. We notice striking dissimilarity across the outputs of the methods, attesting to the difficulty of the task, and to the usefulness of visualizing point tracks as flow maps. We find that when the foreground displacements are large, the flow models often “give up” on the dynamic foreground and produce a motion field that only describes the background. We also find that PIPs++ and LocoTrack often struggle with spatial smoothness, while the flow models do not. CoTracker3 occasionally fails on smoothness too (see row 3 with the car), but less so. Our model appears to produce results that are smooth and accurate, which matches intuitions for a model that blends the 2D processing of flow models with the temporal coherence of point trackers.